Circular No. 7436

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cm IN ANONYMOUS GALAXY J. Mueller reports her discovery of an apparent supernova (mag ∼ 16)

on a near-infrared IVN plate taken on May 31 UT by K. Rykoski with the 1.2-m Oschin Schmidt telescope in the course of the second Palomar Sky Survey. SN 2000cm is located at α = 12^{h}12^{m}59^{s}.1, δ = +7^{o}17′53′′ (equinox 2000.0), which is 9′′ west and 6′′ north of the center of the apparent host galaxy. The new object also appears on other Sky Survey IVN plates taken on June 1 and 3 by Rykoski, but there is no object present at this position on the Digitized Sky Survey or on a second Sky Survey red plate taken on 1996 Apr. 26 by Mueller.

SUPERNOVA 2000cn IN UGC 11064 M. Papenkova and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope of an apparent supernova on unfiltered images taken on June 2.5 (mag ∼ 17.4) and 3.4 UT (mag ∼ 16.9). SN 2000cn is located at α = 17^{h}57^{m}40^{s}.42, δ = +27^{o}49′58′′.1 (equinox 2000.0), which is 7′′.2 west and 7′′.2 south of the nucleus of UGC 11604. An unfiltered image of the field taken on May 29.5 showed nothing at the position of the new object (limiting mag ∼ 19.0).

SUPERNOVA 2000co IN MCG +7-33-17 T. Puckett, Mountain Town, GA, reports the discovery of an appar-

ent supernova (mag 19.2) on an unfiltered CCD frame (limiting mag 20.5) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on June 2.26 UT. SN 2000co is located at α = 15^{h}57^{m}24^{s}.59, δ = +41^{o}54′02′′.0 (equinox 2000.0), which is 4′′.1 east and 9′′.0 south of the center of MCG +7-33-17. The new object was also present on an unfiltered CCD frame taken on June 3.11, but it does not appear on Palomar Sky Survey images taken on 1988 May 15 (limiting mag ∼ 21.5) or 1991 Apr. 10 (limiting mag ∼ 21.5).

CI AQUILAE Visual magnitude estimates: May 8.014 UT, 9.0 (A. Baransky, Kiev,

Ukraine); 16.736, 9.8 (A. Pearce, Nedlands, W. Australia); 21.750, 10.5 (S. Yoshida, Ibaraki, Japan); June 1.842, 10.8 (Pearce).

2000 June 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green